### 13 Dec 2023 v1

# Integral Representations of Three Novel Multiple Zeta Functions for Barnes Type: A Probabilistic Approach

Abstract. Integral representation is one of the powerful tools for studying
analytic continuation of the zeta functions. It is known that Hurwitz zeta
function generalizes the famous Riemann zeta function which plays an
important role in analytic number theory. They both have several
multiple versions in the literature.
In this paper, we introduce three novel multiple
zeta functions for Barnes type and study their integral representations
through hyperbolic probability distributions given by Pitman and Yor
(2003, Canad. J. Math., 55, 292β330). The analytically
continued properties of the three multiple zeta functions are also
investigated. Surprisingly, two of them, unlike the previous
results, can extend analytically
to entire functions in the whole complex plane.
Β
Key words and phrases: Integral representation,
Analytical continuation, Barnes zeta function,
Hyperbolic functions, Hurwitz zeta function, Riemann zeta function.
2020
Mathematics Subject Classifications: Primary 11M06, 11M32, 60E05.
Running title: Integral representations of three novel Barnes
multiple zeta
functions
Postal addresses:
Gwo Dong Lin, Institute of Statistical
Science, Academia Sinica, No.β128, Section 2, Academia Road,
Nankang District, Taipei 11529, Taiwan (ROC). (E-mail:
gdlin@stat.sinica.edu.tw)
Chin-Yuan Hu, National Changhua
University of Education, No.β1, Jinde Road, Changhua 50007, Taiwan
(ROC).
(E-mail: buhuua@gmail.com)
Consider the Hurwitz zeta function
where the parameter and the complex variable has the
real part When , reduces to the
famous Riemann zeta function
which plays an important role in analytic number theory (see, e.g.,
Apostal 1976). It is known that each of both zeta functions in (1)
and (2) can extend to a meromorphic function in the whole complex
plane and has a simple pole at where it has residue 1.
As for multidimensional generalizations of (1) and (2), there are
various multiple (-tuple or -fold) zeta functions proposed in
the literature (see, e.g., the survey paper by Matsumoto 2010, the
thesis by Onozuka 2014, and the references therein), where
is an integer. We recall a few of the results in the following:
(i) EulerβZagier multiple (Riemann) zeta function (Euler 1776,
Zagier 1994, Hoffman 1992, Zhao 2000, Matsumoto 2002a, Mehta et al.
2016)
(ii) Multiple Hurwitz zeta function (Kamano 2006, Murty and Sinha
2006, Kelliher and Masri 2008, Mehta and Viswanadham 2017)
or, in general,
(iii) Barnes multiple zeta function (of one variable) (Barnes 1901, 1904, Choi and Quine 1992, Ruijsenaars 2000, Matsumoto 2002a, 2010)
where and are complex parameters satisfying
some suitable conditions (e.g., all of them have positive real
parts).
For further generalizations of the Barnes multiple zeta function
, see Shintani (1977, 1978) and Mellin (1902, 1904) for
the case of one variable, and Matsumoto (2002a, 2002b, 2003a, 2003b,
2005) for the case of several variables. The analytically continued
properties of the above-mentioned multiple zeta functions have been
investigated. For example, the Barnes multiple zeta function
extends to a meromorphic function in and has
simple poles only at (see Barnes 1904, Ruijsenaars
2000 or Matsumoto 2010). The main tools of analytic continuation of
the zeta functions include integral representation, MellinβBarnes
integral formula, EulerβMaclaurin summation, translation formula
(identity), generalized function, etc. For the recent development on
the multiple zeta functions, see the monograph by Zhao (2016).
In this paper, we will introduce the following three variants (a) β
(c) of the Barnes multiple zeta function, mimicking the
Shintani double zeta functions of the form (Shintani 1975, Kim et
al. 2022)
where denotes the number of distinct solutions to the
quadratic congruence equation
Let be an integer and for convenience let
be
positive real numbers. Denote
and As usual, we
adopt the notations and for real
Then, we define
(a) the multiple hyperbolic sinh-moment function
(b) the multiple hyperbolic cosh-moment function
(c) the multiple hyperbolic tanh-moment function
where are integers,
and
are independently and
identically distributed (i.i.d.) uniform random variables on
It is seen that the function in (3) above
generalizes the Barnes multiple zeta function (see
Corollary 1 below). We will provide the integral representations of
the above three novel multiple zeta functions through the hyperbolic
probability distributions given by Pitman and Yor (2003), and
investigate their analytically continued properties including the
set of singularities. Surprisingly, unlike the previous results, the
functions and in (4) and
(5) above can extend analytically to entire functions in
The main results are stated in Section 3. Section 2 provides the
needed lemmas. Finally, the proofs of the main results are given in
Section 4.
2. Lemmas
We need the following preparatory lemmas. Lemma 1 presents
the basic properties of the gamma function (see, e.g., Apostal 1976,
Chapter 12). Lemma 2 is about Liouvilleβs extension of Dirichlet
integral theorem (see, e.g., Sharma 2005, Chapter 2). The crucial
Lemma 3 can be proved by Moreraβs theorem (see Cheng and Hu 2022).
The hyperbolic random variables in Lemma 4 are introduced and
investigated by Pitman and Yor (2003). On the other hand, for the
derivations of the hyperbolic density functions in Lemma 5, see
Cheng and Hu (2022) as well as Pitman and Yor (2003).
Lemma 1 (Gamma function). Let be the
gamma function defined by
Then the function can extend to a meromorphic function
in the whole complex plane and has simple poles at the points
with residue at where
The gamma function satisfies the equation
It has no zeros, and hence the
reciprocal gamma function, , is an entire function in
with zeros at Moreover, if is
not an integer, then
where denotes the greatest integer less than or equal to
Lemma 2 (LiouvilleβDirichlet integral). Let
be an integer and let be a continuous function on
Then we have
where the parameters and the integral
region
In particular, if where and
then the following formula holds:
Lemma 3 (Integral transform of a distribution). Let be a random variable with distribution on
denoted and let have finite moments of all orders.
For any positive real define the complex-valued function
Then is an entire function in
Remark 1. Note that for
the function value of
becomes a linear combination of moments of Thus, for
simplicity, we might call the integral transformation a
moment function of the distribution
Lemma 4 (Hyperbolic random variables). For
let be three random
variables with characteristic functions
respectively. Then the hyperbolic random variables are infinitely divisible and have finite
moments of all orders.
For
let where are called
hyperbolic -, -, -distributions, respectively.
Lemma 5 (Hyperbolic distributions). For
are absolutely continuous
symmetric distributions about zero and have probability
density functions
respectively, where is the Beta function
In particular, for we have the even density functions:
Throughout the section we assume that is an integer and
that , and
are positive real numbers satisfying Denote
and
Moreover, the
notation ββ means equality in distribution.
Then we have the following results, which as mentioned below
generalize some of the previous results given in Cheng and Hu
(2022).
Theorem 1. Assume that
are independent random variables, where are
hyperbolic sinh random variables defined in Lemma . Then the
multiple hyperbolic sinh-moment function in
has an integral representation of the form:
where
Moreover, the following statements hold true.
The zeta function in can extend
to a meromorphic function in and all the poles are simple.
If the parameter is a positive integer, then the
function in is an analytic function in
except for Precisely,
If the parameter is not an integer, then the zeta
function in has trivial zeros at
and
we have the
following two cases for poles.
For non-negative integer the point is a
simple pole of with residue
For positive integer the point is a
simple pole of with residue
where
If
then and the multiple hyperbolic sinh-moment function
in (3) reduces to the Barnes multiple zeta
function Thus, we have the following.
Corollary 1. The Barnes multiple zeta function
has the integral representation
where
and are independent
random variables, each having the
density function
Moreover, in extends to a meromorphic function in and has simple poles
at
If, further, and then the multiple
hyperbolic sinh-moment function reduces to the
Hurwitz zeta function Thus, we have the following.
Corollary 2 (Cheng and Hu 2022). The Hurwitz
zeta function has the integral representation of the
form
where and
has the density function in
Moreover, in extends to a meromorphic function in and has a simple pole
at
If and then
and the multiple hyperbolic sinh-moment function
in (3) reduces to
Corollary 3 (Cheng and Hu 2022). The function
in has the integral representation
where has the density
function
Moreover, in extends to a meromorphic function in and has simple poles
at
Theorem 2. Assume that
are independent random variables, where are
hyperbolic cosh random variables defined in Lemma Then the
multiple hyperbolic cosh-moment function in
has an integral representation of the form:
where Moreover, the zeta
function in can extend analytically to
an entire function in
If then
and the multiple hyperbolic cosh-moment function
in (4) reduces to the following zeta function
(compare with the zeta function ). Thus, we have the
following.
Corollary 4. The zeta function
in has the integral representation of the form
where
and are
independent random variables, each
having the density function
Moreover, the zeta function in extends to an
entire function in
If, further, and then the zeta
function in (13) reduces to
Thus, we have the following.
Corollary 5 (Cheng and Hu 2022). The zeta
function in has the integral representation
where has the density function in Moreover,
the zeta function in extends to an entire
function in
If and then
and the multiple hyperbolic cosh-moment function
in (4) reduces to
Then we have the following.
Corollary 6 (Cheng and Hu 2022). The zeta
function in has the integral representation
of the form
where has the density
function
Moreover, the zeta function in extends to an
entire function in
Theorem 3. Assume that
are independent random variables, where
are positive integers and are hyperbolic tanh
random variables defined in Lemma Then the multiple hyperbolic
tanh-moment function in has an integral
representation of the form:
where Moreover, the zeta
function in can extend analytically to an
entire function in
If then and
the multiple hyperbolic tanh-moment function in
(5) reduces to
where and
are i.i.d.
uniform random variables on
Thus, we have the
following.
Corollary 7. The multiple hyperbolic
tanh-moment function in
has the integral
representation of the form
where and are i.i.d.
random variables, each having the
density function
Moreover, in extends to an entire function in
If, further, and then the
multiple hyperbolic tanh-moment
function in (21) reduces to the
where has the uniform distribution in
Corollary 8 (Cheng and Hu 2022). The multiple
hyperbolic tanh-moment function in
has the integral representation
of the form
where has the density function
in Moreover, in extends to an entire
function in
If and then
and the multiple hyperbolic tanh-moment function
in (21) reduces to
where are i.i.d. uniform random variables on
Corollary 9 (Cheng and Hu 2022). The zeta
function in has the integral
representation
where has the density
function
Moreover, in extends to an entire function
in
Remark 2. In addition to the above single variable cases, it
is also possible to consider the following cases of several
variables, where
and
in which the parameters and
In these cases, Lemma 2 applies. The results
are too complicated to be presented here.
4. Proofs of the main results
It suffices to prove Theorems 1, 2 and 3, because the corollaries
can be derived easily from the theorems.
Proof of Theorem 1. We first claim that the multiple
series on the RHS of (3) converges absolutely provided To see this, note that by Stirlingβs formula, or more
generally, as
there exist some constants such that
On the other hand, for fixed with take
such that then
Finally, apply the fact that each series
converges.
Multiplying both sides of the
multiple hyperbolic sinh-moment function in (3)
by we have
Recall the gamma identity
and the binomial series
or, equivalently,
Now, by (29), (30), the above claim and Fubiniβs theorem, we can
rewrite (28) as follows:
Next, consider the product term of the integrand in (31) and write,
using Lemma 4,
where
Inserting (32) in (31) yields, by (29) and Fubiniβs theorem again,
where we apply the symmetric property of the hyperbolic
sinh-distribution. This proves the integral representation in (6) by
dividing both sides of (33) by
Recall that the integral transformation is an entire function in
due to Lemma 3. Therefore, the zeta function in
(6) can extend analytically to a meromorphic function in
and all the poles are simple due to Lemma 1. This proves part (a).
Parts (b) and (c) follow from part (a) and Lemma 1 again
immediately. The proof of Theorem 1 is completed.
Proof of Theorem 2. Note first that there is a
relationship between and
specifically, we can rewrite
so that the multiple series on the RHS of (4) also converges
absolutely provided Then proceeding along the
same lines in the proof of Theorem 1, we can prove the theorem as
follows. Multiplying both sides of the multiple hyperbolic
sinh-moment function in (4) by we
have
As before, using the gamma identity and the binomial series, we can
rewrite (34) as follows:
Next, consider the product term of the integrand in (35) and write,
using Lemma 4,
where
Inserting (36) in (35) yields, by (29) and Fubiniβs theorem again,
where we apply the symmetric property of the hyperbolic
cosh-distribution. By cancelling the common term in both
sides of (37), we
proves the integral representation in (9).
Recall again that the integral transformation is an entire function in due to
Lemma 3. Therefore, the zeta function in (12)
can extend analytically to an entire function in This
completes the proof of Theorem 2.
Proof of Theorem 3. As in the proof of Theorem 2,
the multiple series on the RHS of (5) converges absolutely provided
by noting that
Multiplying both sides of the function in (5)
by we have
Then, by (29) and Fubiniβs theorem, we can rewrite the expectation
where we use the fact that for uniform random variable on
Next, recall that
and that
Then it follows from (38)β(41) and the definition of the random
variable that
Cancelling in both sides of (42), we obtain the integral
representation in (20). Finally, Lemma 3 completes the proof of the
theorem.

### 02 Jan 2024 v2

@@
 be
-positive real numbers. Denote
-and As usual, we
-adopt the notations and for real
+positive real numbers. Denote and
+As usual, we adopt the
+notations and for real
 Then, we define
@@
 identically distributed (i.i.d.) uniform random variables on
-It is seen that the function in (3) above
-generalizes the Barnes multiple zeta function (see
-Corollary 1 below). We will provide the integral representations of
-the above three novel multiple zeta functions through the hyperbolic
-probability distributions given by Pitman and Yor (2003), and
-investigate their analytically continued properties including the
-set of singularities. Surprisingly, unlike the previous results, the
-functions and in (4) and
-(5) above can extend analytically to entire functions in
-The main results are stated in Section 3. Section 2 provides the
-needed lemmas. Finally, the proofs of the main results are given in
-Section 4.
+It is seen that when all the function
+in (3) above reduces to the Barnes
+multiple zeta function (see Corollary 1 below). We will
+provide the integral representations of the above three novel
+multiple zeta functions through the hyperbolic probability
+distributions given by Pitman and Yor (2003), and investigate their
+analytically continued properties including the set of
+singularities. Surprisingly, unlike the previous results, the
+functions and in (4) and (5) above can extend analytically to
+entire functions in The main results are stated in Section
+3. Section 2 provides the needed lemmas. Finally, the proofs of the
+main results are given in Section 4.
 2. Lemmas
@@
 Chapter 12). Lemma 2 is about Liouvilleβs extension of Dirichlet
-integral theorem (see, e.g., Sharma 2005, Chapter 2). The crucial
-Lemma 3 can be proved by Moreraβs theorem (see Cheng and Hu 2022).
-The hyperbolic random variables in Lemma 4 are introduced and
+integral theorem (see, e.g., Sharma 2005, Chapter 2, or Gibson
+1931, p.β492). The crucial Lemma 3 modifies slightly
+Cheng and Huβs (2022) Lemma 2.1. The
+hyperbolic random variables in Lemma 4 are introduced and
 investigated by Pitman and Yor (2003). On the other hand, for the
@@
 Lemma 2 (LiouvilleβDirichlet integral). Let
-be an integer and let be a continuous function on
-Then we have
-where the parameters and the integral
-region
+be an integer and let be a Lebesgue measurable function
+on Then we have
+provided that the resulting single integral on the right exists.
+Here, the parameters and the region of
+multiple integral
 In particular, if where and
@@
 denoted and let have finite moments of all orders.
-For any positive real define the complex-valued function
-Then is an entire function in
-Remark 1. Note that for
-the function value of
-becomes a linear combination of moments of Thus, for
-simplicity, we might call the integral transformation a
-moment function of the distribution
+For any nonzero real define the complex-valued function
+where Then is an
+entire function in
+Proof. We sketch the proof below. For each define the
+closed square
+Denote the integer Then there exists a constant
+such that
+This implies that the integral transformation is
+absolutely convergent and converges uniformly in each compact
+subset of Moreover, is a continuous function in
+Let be any closed triangle in and recall
+that for each fixed real is an entire function in
+Then we have, by Cauchyβs theorem,
+where is the boundary of Finally,
+Moreraβs theorem completes the proof.
+Remark 1. Note that for the
+function value of becomes a linear combination of moments
+of Thus, for simplicity, we might call the integral
+transformation a moment function of the distribution
 Lemma 4 (Hyperbolic random variables). For
@@
 Throughout the section we assume that is an integer and
-that , and
-are positive real numbers satisfying Denote
+that , and are
+positive real numbers satisfying Denote
 and
@@
 hyperbolic sinh random variables defined in Lemma . Then the
-multiple hyperbolic sinh-moment function in
-has an integral representation of the form:
+multiple hyperbolic sinh-moment function in has an integral representation of the form:
 where
 Moreover, the following statements hold true.
-The zeta function in can extend
+The zeta function in
+can extend
 to a meromorphic function in and all the poles are simple.
 If the parameter is a positive integer, then the
-function in is an analytic function in
-except for Precisely,
+function in is an analytic
+function in except for Precisely,
 If the parameter is not an integer, then the zeta
-function in has trivial zeros at
-and
+function in has trivial zeros
+at and
 we have the
@@
 where
-If
-then and the multiple hyperbolic sinh-moment function
-in (3) reduces to the Barnes multiple zeta
-function Thus, we have the following.
+If then
+and the multiple hyperbolic sinh-moment function in (3) reduces to the Barnes multiple zeta function
+Thus, we have the following.
 Corollary 1. The Barnes multiple zeta function
 has the integral representation
-where
-and are independent
-random variables, each having the
-density function
+where and are
+independent random variables, each
+having the density function
 Moreover, in extends to a meromorphic function in and has simple poles
@@
 If, further, and then the multiple
-hyperbolic sinh-moment function reduces to the
-Hurwitz zeta function Thus, we have the following.
+hyperbolic sinh-moment function
+reduces to the Hurwitz zeta function Thus, we have the
+following.
 Corollary 2 (Cheng and Hu 2022). The Hurwitz
@@
 If and then
-and the multiple hyperbolic sinh-moment function
-in (3) reduces to
+and the multiple hyperbolic sinh-moment function in (3) reduces to
 Corollary 3 (Cheng and Hu 2022). The function
 in has the integral representation
-where has the density
-function
+where and has
+the density function
 Moreover, in extends to a meromorphic function in and has simple poles
@@
 hyperbolic cosh random variables defined in Lemma Then the
-multiple hyperbolic cosh-moment function in
-has an integral representation of the form:
+multiple hyperbolic cosh-moment function in has an integral representation of the form:
 where Moreover, the zeta
-function in can extend analytically to
+function in can extend
+analytically to
 an entire function in
 If then
-and the multiple hyperbolic cosh-moment function
-in (4) reduces to the following zeta function
+and the multiple hyperbolic cosh-moment function in (4) reduces to the following zeta function
 (compare with the zeta function ). Thus, we have the
@@
 in has the integral representation of the form
-where
-and are
+where and are
 independent random variables, each
@@
 function in has the integral representation
-where has the density function in Moreover,
-the zeta function in extends to an entire
-function in
+where and has the density function
+in Moreover, the zeta function in
+extends to an entire function in
 If and then
-and the multiple hyperbolic cosh-moment function
-in (4) reduces to
+and the multiple hyperbolic cosh-moment function in (4) reduces to
 Then we have the following.
@@
 of the form
-where has the density
-function
+where and has
+the density function
 Moreover, the zeta function in extends to an
@@
 random variables defined in Lemma Then the multiple hyperbolic
-tanh-moment function in has an integral
-representation of the form:
+tanh-moment function in has
+an integral representation of the form:
 where Moreover, the zeta
-function in can extend analytically to an
-entire function in
+function in can extend
+analytically to an entire function in
 If then and
-the multiple hyperbolic tanh-moment function in
-(5) reduces to
+the multiple hyperbolic tanh-moment function in (5) reduces to
 where and
@@
 multiple hyperbolic tanh-moment
-function in (21) reduces to the
+function in (21) reduces to
 where has the uniform distribution in
@@
 in which the parameters and
-In these cases, Lemma 2 applies. The results
-are too complicated to be presented here.
+In these cases,
+Lemma 2 applies. The results are too complicated to be presented
+here.
 4. Proofs of the main results
@@
 Multiplying both sides of the
-multiple hyperbolic sinh-moment function in (3)
-by we have
+multiple hyperbolic sinh-moment function in (3) by we have
 Recall the gamma identity
-and the binomial series
-or, equivalently,
+and the binomial series or, equivalently,
 Now, by (29), (30), the above claim and Fubiniβs theorem, we can
@@
 Recall that the integral transformation is an entire function in
-due to Lemma 3. Therefore, the zeta function in
-(6) can extend analytically to a meromorphic function in
-and all the poles are simple due to Lemma 1. This proves part (a).
-Parts (b) and (c) follow from part (a) and Lemma 1 again
+due to Lemma 3. Therefore, the zeta function in (6) can extend analytically to a meromorphic function
+in and all the poles are simple due to Lemma 1. This proves
+part (a). Parts (b) and (c) follow from part (a) and Lemma 1 again
 immediately. The proof of Theorem 1 is completed.
 Proof of Theorem 2. Note first that there is a
-relationship between and
-specifically, we can rewrite
+relationship between and specifically, we can rewrite
 so that the multiple series on the RHS of (4) also converges
@@
 follows. Multiplying both sides of the multiple hyperbolic
-sinh-moment function in (4) by we
-have
+sinh-moment function in (4) by
+we have
 As before, using the gamma identity and the binomial series, we can
@@
 Recall again that the integral transformation is an entire function in due to
-Lemma 3. Therefore, the zeta function in (12)
-can extend analytically to an entire function in This
-completes the proof of Theorem 2.
+Lemma 3. Therefore, the zeta function in (12) can extend analytically to an entire function in
+This completes the proof of Theorem 2.
 Proof of Theorem 3. As in the proof of Theorem 2,
@@
 by noting that
-Multiplying both sides of the function in (5)
-by we have
+Multiplying both sides of the function in (5) by we have
 Then, by (29) and Fubiniβs theorem, we can rewrite the expectation

### 31 Mar 2026 v3

@@
 through hyperbolic probability distributions given by Pitman and Yor
-(2003, Canad. J. Math., 55, 292β330). The analytically
+(Canad. J. Math., 55 (2003), 292–330). The analytically
 continued properties of the three multiple zeta functions are also
-investigated. Surprisingly, two of them, unlike the previous
-results, can extend analytically
-to entire functions in the whole complex plane.
-Β
+investigated. The first one generalizes the well-known Barnes
+multiple zeta function, while the second and the third, unlike the
+previous results, can extend analytically to entire functions in the
+whole complex plane. These are somewhat surprising results.
 Key words and phrases: Integral representation,
@@
 Mathematics Subject Classifications: Primary 11M06, 11M32, 60E05.
-Running title: Integral representations of three novel Barnes
+Running title: Three novel Barnes
 multiple zeta
@@
 Gwo Dong Lin, Institute of Statistical
-Science, Academia Sinica, No.β128, Section 2, Academia Road,
+Science, Academia Sinica, No. 128, Section 2, Academia Road,
 Nankang District, Taipei 11529, Taiwan (ROC). (E-mail:
@@
 Chin-Yuan Hu, National Changhua
-University of Education, No.β1, Jinde Road, Changhua 50007, Taiwan
+University of Education, No. 1, Jinde Road, Changhua 50007, Taiwan
 (ROC).
@@
 famous Riemann zeta function
-which plays an important role in analytic number theory (see, e.g.,
-Apostal 1976). It is known that each of both zeta functions in (1)
-and (2) can extend to a meromorphic function in the whole complex
-plane and has a simple pole at where it has residue 1.
+which plays an important role in analytic number theory (see, e.g., Apostal
+[1]).
+It is known that each of both zeta functions in (1) and (2) can
+extend to a meromorphic function in the whole complex plane and
+has a simple pole at where it has residue 1.
 As for multidimensional generalizations of (1) and (2), there are
 various multiple (-tuple or -fold) zeta functions proposed in
-the literature (see, e.g., the survey paper by Matsumoto 2010, the
-thesis by Onozuka 2014, and the references therein), where
+the literature (see, e.g., the survey paper by Matsumoto
+[17], the thesis by Onozuka [23], and the
+references therein), where
 is an integer. We recall a few of the results in the following:
-(i) EulerβZagier multiple (Riemann) zeta function (Euler 1776,
-Zagier 1994, Hoffman 1992, Zhao 2000, Matsumoto 2002a, Mehta et al.
-2016)
-(ii) Multiple Hurwitz zeta function (Kamano 2006, Murty and Sinha
-2006, Kelliher and Masri 2008, Mehta and Viswanadham 2017)
+(i) Euler–Zagier multiple (Riemann) zeta function ([6, 8, 12, 18, 30, 31])
+(ii) Multiple Hurwitz zeta function ([9, 10, 19, 22])
 or, in general,
-(iii) Barnes multiple zeta function (of one variable) (Barnes 1901, 1904, Choi and Quine 1992, Ruijsenaars 2000, Matsumoto 2002a, 2010)
+(iii) Barnes multiple zeta function (of one variable) ([2, 3, 5, 12, 17, 25])
 where and are complex parameters satisfying
@@
 For further generalizations of the Barnes multiple zeta function
-, see Shintani (1977, 1978) and Mellin (1902, 1904) for
-the case of one variable, and Matsumoto (2002a, 2002b, 2003a, 2003b,
-2005) for the case of several variables. The analytically continued
-properties of the above-mentioned multiple zeta functions have been
-investigated. For example, the Barnes multiple zeta function
-extends to a meromorphic function in and has
-simple poles only at (see Barnes 1904, Ruijsenaars
-2000 or Matsumoto 2010). The main tools of analytic continuation of
-the zeta functions include integral representation, MellinβBarnes
-integral formula, EulerβMaclaurin summation, translation formula
-(identity), generalized function, etc. For the recent development on
-the multiple zeta functions, see the monograph by Zhao (2016).
-In this paper, we will introduce the following three variants (a) β
+, see [20, 21, 28, 29] for the
+case of one variable, and [12]–[16] for the
+case of several variables. The analytically continued properties of
+the above-mentioned multiple zeta functions have been investigated.
+For example, the Barnes multiple zeta function extends
+to a meromorphic function in and has simple poles only at
+(see [3, 17, 25]). The main
+tools of analytic continuation of the zeta functions include
+integral representation, Mellin–Barnes integral formula,
+Euler–Maclaurin summation, translation formula (identity),
+generalized function, etc. For the recent development on the
+multiple zeta functions, see the monograph [32].
+In this paper, we will introduce the following three variants (a) –
 (c) of the Barnes multiple zeta function, mimicking the
-Shintani double zeta functions of the form (Shintani 1975, Kim et
-al. 2022)
+Shintani double zeta functions of the form ([11, 27])
 where denotes the number of distinct solutions to the
@@
 notations and for real
-Then, we define
+Then, we define (see Remark 1 below for the names of moment functions)
 (a) the multiple hyperbolic sinh-moment function
@@
 multiple zeta functions through the hyperbolic probability
-distributions given by Pitman and Yor (2003), and investigate their
-analytically continued properties including the set of
+distributions given by Pitman and Yor [24], and investigate
+their analytically continued properties including the set of
 singularities. Surprisingly, unlike the previous results, the
@@
 main results are given in Section 4.
-2. Lemmas
 We need the following preparatory lemmas. Lemma 1 presents
-the basic properties of the gamma function (see, e.g., Apostal 1976,
-Chapter 12). Lemma 2 is about Liouvilleβs extension of Dirichlet
-integral theorem (see, e.g., Sharma 2005, Chapter 2, or Gibson
-1931, p.β492). The crucial Lemma 3 modifies slightly
-Cheng and Huβs (2022) Lemma 2.1. The
+the basic properties of the gamma function (see, e.g., Apostal [1, Chapter
+12]). Lemma 2 is about Liouville’s extension of Dirichlet
+integral theorem (see, e.g., Sharma [26, Chapter 2], or Gibson
+[7, page 492]). The crucial Lemma 3 modifies slightly
+Cheng and Hu’s [4] Lemma 2.1. The
 hyperbolic random variables in Lemma 4 are introduced and
-investigated by Pitman and Yor (2003). On the other hand, for the
-derivations of the hyperbolic density functions in Lemma 5, see
-Cheng and Hu (2022) as well as Pitman and Yor (2003).
+investigated by Pitman and Yor [24]. On the other hand, for
+the derivations of the hyperbolic density functions in Lemma 5, see
+[4, 24].
 Lemma 1 (Gamma function). Let be the
@@
 where denotes the greatest integer less than or equal to
-Lemma 2 (LiouvilleβDirichlet integral). Let
+Lemma 2 (Liouville–Dirichlet integral). Let
 be an integer and let be a Lebesgue measurable function
@@
 In particular, if where and
-then the following formula holds:
+then
 Lemma 3 (Integral transform of a distribution). Let be a random variable with distribution on
@@
 that for each fixed real is an entire function in
-Then we have, by Cauchyβs theorem,
+Then we have, by applying first Fubini’s theorem and then
+Cauchy’s theorem,
 where is the boundary of Finally,
-Moreraβs theorem completes the proof.
+Morera’s theorem completes the proof.
 Remark 1. Note that for the
@@
 Moreover, the
-notation ββ means equality in distribution.
+notation “” means equality in distribution.
 Then we have the following results, which as mentioned below
 generalize some of the previous results given in Cheng and Hu
-(2022).
+[4].
 Theorem 1. Assume that
@@
 following.
-Corollary 2 (Cheng and Hu 2022). The Hurwitz
-zeta function has the integral representation of the
-form
+Corollary 2 ([4]). The Hurwitz zeta
+function has the integral representation
 where and
@@
 and the multiple hyperbolic sinh-moment function in (3) reduces to
-Corollary 3 (Cheng and Hu 2022). The function
+Corollary 3 ([4]). The function
 in has the integral representation
@@
 Corollary 4. The zeta function
-in has the integral representation of the form
+in has the integral representation
 where and are
@@
 Thus, we have the following.
-Corollary 5 (Cheng and Hu 2022). The zeta
-function in has the integral representation
+Corollary 5 ([4]). The zeta function
+in has the integral representation
 where and has the density function
@@
 Then we have the following.
-Corollary 6 (Cheng and Hu 2022). The zeta
-function in has the integral representation
-of the form
+Corollary 6 ([4]). The zeta function
+in has the integral representation
 where and has
@@
 has the integral
-representation of the form
+representation
 where and are i.i.d.
@@
 where has the uniform distribution in
-Corollary 8 (Cheng and Hu 2022). The multiple
+Corollary 8 ([4]). The multiple
 hyperbolic tanh-moment function in
 has the integral representation
-of the form
 where has the density function
@@
 where are i.i.d. uniform random variables on
-Corollary 9 (Cheng and Hu 2022). The zeta
-function in has the integral
-representation
+Corollary 9 ([4]). The zeta function
+in has the integral representation
 where has the density
@@
 Proof of Theorem 1. We first claim that the multiple
-series on the RHS of (3) converges absolutely provided To see this, note that by Stirlingβs formula, or more
+series on the RHS of (3) converges absolutely provided To see this, note that by Stirling’s formula, or more
 generally, as
@@
 and the binomial series or, equivalently,
-Now, by (29), (30), the above claim and Fubiniβs theorem, we can
+Now, by (29), (30), the above claim and Fubini’s theorem, we can
 rewrite (28) as follows:
@@
 where
-Inserting (32) in (31) yields, by (29) and Fubiniβs theorem again,
+Inserting (32) in (31) yields, by (29) and Fubini’s theorem again,
 where we apply the symmetric property of the hyperbolic
@@
 where
-Inserting (36) in (35) yields, by (29) and Fubiniβs theorem again,
+Inserting (36) in (35) yields, by (29) and Fubini’s theorem again,
 where we apply the symmetric property of the hyperbolic
@@
 Multiplying both sides of the function in (5) by we have
-Then, by (29) and Fubiniβs theorem, we can rewrite the expectation
+Then, by (29) and Fubini’s theorem, we can rewrite the expectation
 where we use the fact that for uniform random variable on
@@
 and that
-Then it follows from (38)β(41) and the definition of the random
+Then it follows from (38)–(40) and the definition of the random
 variable that